40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Wed May 12 15:43:30 2021, Aidan, Summary, 2um Photodiodes, Chamber is leaking 
    Reply  Wed May 26 18:38:25 2021, Aidan, Summary, 2um Photodiodes, Chamber is leaking 
       Reply  Fri Jun 11 07:42:54 2021, Stephen, Summary, 2um Photodiodes, Chamber is leaking IMG_8865.JPGIMG_8881.JPG
          Reply  Fri Jun 11 16:48:31 2021, Aidan, Stephen, Summary, 2um Photodiodes, Chamber is leaking IMG_3002.jpgIMG_3001.jpg
Message ID: 2588     Entry time: Fri Jun 11 16:48:31 2021     In reply to: 2586
Author: Aidan, Stephen 
Type: Summary 
Category: 2um Photodiodes 
Subject: Chamber is leaking 

We hit another dead-end with leak hunting the IR labs dewer (we replaced screws and helicoil on the valve connection but there is still a big leak). We cleaned the flange and O-ring with isopropanal and replaced the threads with helicoil but still get the same sort of leak where we only hit 1E-2 Torr after 5 minutes of pumping and stablize around 1E-3.

After turning off the pumping station, the pressure rose quickly to 1Torr (in roughly 10 minutes or so).

 

Quote:

[Stephen, Aidan, Wednesday 09 June]

Summary and Plan:

  • Poor sealing at the output valve (The Gap) needs to be resolved.
  • Planning to install #4-40 helicoils today (chamber will remain sealed, will need to remove output valve and cover output orifice, then transport the chamber to the WB EE shop for redrilling of holes.)
  • Meeting with Nina and Aidan this afternoon to iterate one more time.

Troubleshooting steps taken:

  1. Aidan took us through the full sequence of pump down and disassembly to bring me up to speed.
  2. We opened the lid and inspected the old o-ring.
    1. Signs of plastic deformation and of small flecks of particulate near sealing surface - good idea to change.
  3. We found new o-rings in a box from the Cryo lab, and one of these was swapped in after a good wipedown with IPA.
  4. Upon pumpdown, Aidan compared behavior and found no meaningful change to rate of pumpdown or stable pressure in e-4 torr range after 10+ minutes.
    1. By valving off [chamber + gauge] from pump line, it was clear that there was a leak in that volume, as within seconds the pressure rose from e-4 torr to e-2 torr, and stabilized at _(need to confirm - e0?)_ torr over ~10 minutes.
  5. Attempted to squirt IPA along o-ring seals, but there was not good access to the sealing surfaces, so this was a null test
  6. Looked closer at all of the chamber features, and noticed The Gap between chamber wall and chamber output valve, pictured in [Attachment 1]. Not good! But promising as a leak source.
    1. Three of the four screws were found to be loose due to apparent thread damage.
  7. IPA was squirted into The Gap at stable pressure of e-4 torr, but no change in pressure was noticed.
  8. Longer #4-40 screw reinstallation was attempted, and I could feel a small amount of pull at the very tip of the screw, but tightening the screws led to that small pull to fail as well - need to rework.
  9. The Clamp was installed and The Gap was closed [Attachment 2].
    1. When isolated from the vacuum pump, the chamber pressure progressed more slowly. Within seconds, we were at e-3 torr, and over 10 minutes the pressure stabled at e-2 torr, about 30x lower pressure per Aidan's records.

 

 
Attachment 1: IMG_3002.jpg  2.595 MB  Uploaded Fri Jun 11 17:48:42 2021  | Hide | Hide all
IMG_3002.jpg
Attachment 2: IMG_3001.jpg  2.083 MB  Uploaded Fri Jun 11 17:48:51 2021  | Hide | Hide all
IMG_3001.jpg
ELOG V3.1.3-